I answered the equation 9x^2=36 in what i thought was quadratic form but my professor informed me that it was not, i got the answer 2,-2 which was correct.. what are the correct steps to this problem?

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

9x^2=36
9x^2-36=0=========THIS IS YOUR QUADRATIC
9x^2-36=0
0=0
solution 2,-2

Answer by Alan3354(69443)

I answered the equation 9x^2=36 in what i thought was quadratic form but my professor informed me that it was not, i got the answer 2,-2 which was correct.. what are the correct steps to this problem?
-------------
9x^2 = 36
x^2 = 4
x = -2, +2
---------
Or
9x^2 - 36 = 0
x^2 - 4 = 0
(x-2)*(x+2) = 0
same answer